# Improving GFlowNets for Text-to-Image Diffusion Alignment

### v1

@@ 1 Introduction
Diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020) have drawn significant attention in machine learning due to their impressive capability to generate high-quality visual data and applicability across a diverse range of domains, including text-to-image synthesis (Rombach et al., 2021), 3D generation (Poole et al., 2022), material design (Yang et al., 2023), protein conformation modeling (Abramson et al., 2024), and continuous control (Janner et al., 2022). These models, through a process of gradually denoising a random distribution, learn to replicate complex data distributions, showcasing their robustness and flexibility. The traditional training of diffusion models typically relies on large datasets, from which the models learn to generate new samples that mimic and interpolate the observed examples.
However, such a dataset-dependent approach often overlooks the opportunity to control and direct the generation process towards outputs that not only resemble the training data but also possess specific, desirable properties (Lee et al., 2023). These properties are often defined through explicit reward functions that assess certain properties, such as the aesthetic quality of images. Such a requirement is crucial in fields where adherence to particular characteristics is necessary, such as alignment or drug discovery. The need to integrate explicit guidance without relying solely on datasets presents a unique challenge for training methodologies. Previous works have utilized methods such as reinforcement learning (RL) (Black et al., 2023; Fan et al., 2023) to tackle this problem. Nonetheless, these methods still suffer from issues like low sample efficiency.
In this work, we propose a novel approach, diffusion alignment with GFlowNets (DAG), that fine-tunes diffusion models to optimize black-box reward functions directly. Generative flow networks (Bengio et al., 2023, GFlowNets), initially introduced for efficient probabilistic inference with given densities in structured spaces, provide a unique framework for this task. Though initially proposed for composite graph-like structures, prior works have extended the GFlowNet framework to diffusion modeling (Zhang et al., 2022a; Lahlou et al., 2023). This work further investigates GFlowNet-inspired algorithms for the task of text-to-image diffusion alignment. By aligning the learning process to focus on generating samples with probability proportional to reward functions rather than maximizing them, our method allows the diffusion model to directly target and generate samples that are not only high in quality but also fulfill specific predefined criteria. Besides developing a denoising diffusion probabilistic model-specific GFlowNet algorithm, we also propose a new KL-based way to optimize our models. In summary, our contributions are as follows:
@@
2 Preliminaries
2.1 Diffusion models
@@
Denoising diffusion model (Vincent, 2011; Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2020) is a class of hierarchical latent variable models. The latent variables are initialized from a white noise $\mathbf{x}_{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ and then go through a sequential denoising (reverse) process $p_{\boldsymbol{\theta}}(\mathbf{x}_{t-1}|\mathbf{x}_{t})$. Therefore, the resulting generated distribution takes the form of
@@
On the other hand, the variational posterior $q(\mathbf{x}_{1:T}|\mathbf{x}_{0})$, also called a diffusion or forward process, can be factorized as a Markov chain $\prod_{t=1}^{T}q(\mathbf{x}_{t}|\mathbf{x}_{t-1})$ composed by a series of conditional Gaussian distributions $q(\mathbf{x}_{t}|\mathbf{x}_{t-1})=\mathcal{N}(\mathbf{x}_{t};{\alpha_{t}/\alpha_{t-1}}\mathbf{x}_{t-1},(1-\alpha_{t}^{2}/\alpha_{t-1}^{2})\mathbf{I})$, where $\{\alpha_{t},\sigma_{t}\}_{t}$ is a set of pre-defined signal-noise schedule. Specifically, in Ho et al. (2020) we have $\alpha_{t}^{2}+\sigma_{t}^{2}=1$. The benefit of such a noising process is that its marginal has a simple close form: $q(\mathbf{x}_{t}|\mathbf{x}_{0})=\int q(\mathbf{x}_{1:t}|\mathbf{x}_{0})\mathop{}\!\mathrm{d}\mathbf{x}_{1:t-1}=\mathcal{N}(\mathbf{x}_{t};\alpha_{t}\mathbf{x}_{0},\sigma_{t}^{2}\mathbf{I})$.
@@
where $\hat{\mathbf{x}}_{\boldsymbol{\theta}}(\mathbf{x}_{t},t)$ is a deep neural network to predict the original clean data $\mathbf{x}_{0}$ given the noisy input $\mathbf{x}_{t}=\alpha_{t}\mathbf{x}_{0}+\sigma_{t}{\boldsymbol{\epsilon}}$, which can be used to parameterize the denoising process $p_{\boldsymbol{\theta}}(\mathbf{x}_{t-1}|\mathbf{x}_{t})=\mathcal{N}(\mathbf{x}_{t-1};\left(\sigma_{t-1}^{2}\alpha_{t}\mathbf{x}_{t}+(\alpha_{t-1}^{2}-\alpha_{t}^{2})\hat{\mathbf{x}}_{\boldsymbol{\theta}}(\mathbf{x}_{t},t)\right)/\sigma_{t}^{2}\alpha_{t-1},(1-\alpha_{t}^{2}/\alpha_{t-1}^{2})\mathbf{I})$. In practice, the network can also be parameterized with noise prediction or v-prediction (Salimans & Ho, 2022). The network architecture usually has a U-Net (Ronneberger et al., 2015) structure.
@@
2.2 GFlowNets
Generative flow network (Bengio et al., 2021, GFlowNet) is a high-level algorithmic framework of amortized inference, also known as training generative models with a given unnormalized target density function. Let $\mathcal{G}=(\mathcal{S},\mathcal{A})$ be a directed acyclic graph, where $\mathcal{S}$ is the set of states and $\mathcal{A}\subseteq\mathcal{S}\times\mathcal{S}$ are the set of actions. We assume the environmental transition is deterministic, i.e., one action would only lead to one next state. There is a unique initial state $\mathbf{s}_{0}\in\mathcal{S}$ which has no incoming edges and a set of terminal states $\mathbf{s}_{N}$ without outgoing edges. A GFlowNet has a stochastic forward policy $P_{F}(\mathbf{s}^{\prime}|\mathbf{s})$ for transition $(\mathbf{s}\to\mathbf{s}^{\prime})$ as a conditional distribution over the children of a given state $\mathbf{s}$, which can be used to induce a distribution over trajectories via $P(\tau)=\prod_{n=0}^{N-1}P_{F}(\mathbf{s}_{n+1}|\mathbf{s}_{n})$, where $\tau=(\mathbf{s}_{0},\mathbf{s}_{1},\ldots,\mathbf{s}_{N})$. On the other hand, the backward policy $P_{B}(\mathbf{s}|\mathbf{s}^{\prime})$ is a distribution over the parents of a given state $\mathbf{s}^{\prime}$. The terminating distribution defined by $P_{T}(\mathbf{x})=\sum_{\tau\to\mathbf{x}}P_{F}(\tau)$ is the ultimate terminal state distribution generated by the GFlowNet. The goal of training GFlowNet is to obtain a forward policy such that $P_{T}(\cdot)\propto R(\cdot)$, where $R(\cdot)$ is a black-box reward function or unnormalized density that takes only non-negative values. Notice that we do not know the normalizing factor $Z=\sum_{\mathbf{x}}R(\mathbf{x})$. We can use the trajectory flow function $F(\tau)=ZP_{F}(\tau)$ to take in the effect of the normalizing factor, and the corresponding state flow function $F(\mathbf{s})=\sum_{\tau\ni\mathbf{s}}F(\tau)$ to model the unnormalized probability flow of intermediate state $\mathbf{s}$.

### v2

@@ 1 Introduction
Diffusion models (SohlDickstein2015DeepUL; Ho2020DenoisingDP) have drawn significant attention in machine learning due to their impressive capability to generate high-quality visual data and applicability across a diverse range of domains, including text-to-image synthesis (Rombach2021HighResolutionIS), 3D generation (poole2022dreamfusion), material design (yang2023scalable), protein conformation modeling (abramson2024accurate), and continuous control (janner2022planning). These models, through a process of gradually denoising a random distribution, learn to replicate complex data distributions, showcasing their robustness and flexibility. The traditional training of diffusion models typically relies on large datasets, from which the models learn to generate new samples that mimic and interpolate the observed examples.
However, such a dataset-dependent approach often overlooks the opportunity to control and direct the generation process towards outputs that not only resemble the training data but also possess specific, desirable properties (Lee2023AligningTM). These properties are often defined through explicit reward functions that assess certain properties, such as the aesthetic quality of images. Such a requirement is crucial in fields where adherence to particular characteristics is necessary, such as alignment or drug discovery. The need to integrate explicit guidance without relying solely on datasets presents a unique challenge for training methodologies. Previous works have utilized methods such as reinforcement learning (RL) (Black2023TrainingDM; Fan2023DPOKRL) to tackle this problem. Nonetheless, these methods still suffer from issues like low sample efficiency.
In this work, we propose a novel approach, diffusion alignment with GFlowNets (DAG), that fine-tunes diffusion models to optimize black-box reward functions directly. Generative flow networks (bengio2021foundations, GFlowNets), initially introduced for efficient probabilistic inference with given densities in structured spaces, provide a unique framework for this task. Though initially proposed for composite graph-like structures, prior works have extended the GFlowNet framework to diffusion modeling (zhang2022unifying; lahlou2023cgfn). This work further investigates GFlowNet-inspired algorithms for the task of text-to-image diffusion alignment. By aligning the learning process to focus on generating samples with probability proportional to reward functions rather than maximizing them, our method allows the diffusion model to directly target and generate samples that are not only high in quality but also fulfill specific predefined criteria. Besides developing a denoising diffusion probabilistic model-specific GFlowNet algorithm, we also propose a new KL-based way to optimize our models. In summary, our contributions are as follows:
@@
2 Methodology
2.1 Diffusion alignment with GFlowNets
We review the recipe about viewing the denoising process as a MDP in Section D.1. In this section, we describe our proposed algorithm, diffusion alignment with GFlowNets (DAG). Rather than directly optimizing the reward targets as in RL, we aim to train the generative models so that in the end they could generate objects with a probability proportional to the reward function: $p_{\boldsymbol{\theta}}(\mathbf{x}_{0})\propto R(\mathbf{x}_{0})$. To achieve this, we construct the following DB-based training objective based on Equation 14, by regressing its one side to another in the logarithm scale for any diffusion step transition $(\mathbf{x}_{t},\mathbf{x}_{t-1})$.
$\ell_{\text{DB}}(\mathbf{x}_{t},$ $\mathbf{x}_{t-1})=(\log F_{\boldsymbol{\phi}}(\mathbf{x}_{t},t)+\log p_{\boldsymbol{\theta}}(\mathbf{x}_{t-1}|\mathbf{x}_{t},t)$ (1)
$-\log F_{\boldsymbol{\phi}}(\mathbf{x}_{t-1},t-1)-\log q(\mathbf{x}_{t}|\mathbf{x}_{t-1}))^{2}$ (2)
We additionally force $F_{\boldsymbol{\phi}}(\mathbf{x}_{t},t=0)=R(\mathbf{x}_{0})$ to introduce the reward signal. Here ${\boldsymbol{\theta}},{\boldsymbol{\phi}}$ are the parameters of the diffusion U-Net model and the GFlowNet state flow function (which is another neural network), respectively. One can prove that if the optimization is perfect, the resulting model will generate a distribution whose density value is proportional to the reward function $R(\cdot)$ (bengio2021foundations; Zhang2023DiffusionGF).
Algorithm 1 Diffusion alignment with GFlowNets (DAG-DB & DAG-KL)
0:  Denoising policy $p_{\boldsymbol{\theta}}(\mathbf{x}_{t-1}|\mathbf{x}_{t},t)$, noising policy $q(\mathbf{x}_{t}|\mathbf{x}_{t-1})$, flow function $F_{\boldsymbol{\phi}}(\mathbf{x}_{t},t)$, black-box reward function $R(\cdot)$
1:  repeat
2:     Rollout $\tau=\{\mathbf{x}_{t}\}_{t}$ with $p_{\boldsymbol{\theta}}(\mathbf{x}_{t-1}|\mathbf{x}_{t},t)$
3:     For each transition $(\mathbf{x}_{t},\mathbf{x}_{t-1})\in\tau$:
4:     if algorithm is DAG-DB then
5:         # normal DB-based update
6:        Update ${\boldsymbol{\theta}}$ and ${\boldsymbol{\phi}}$ with Equation 5
7:     else if algorithm is DAG-KL then
8:         # KL-based update
9:        Update ${\boldsymbol{\phi}}$ with Equation 5
10:        Update ${\boldsymbol{\theta}}$ with Equation 11
11:     end if
12:  until some convergence condition
One way to parameterize the state flow function $F$ is through the so-called forward-looking (pan2023better, FL) technique in the way of $F_{\boldsymbol{\phi}}(\mathbf{x}_{t},t)=\tilde{F}_{\boldsymbol{\phi}}(\mathbf{x}_{t},t)R(\mathbf{x}_{t})$, where $\tilde{F}_{\boldsymbol{\phi}}$ is the actual neural network to be learned. Intuitively, this is equivalent to initializing the state flow function to be the reward function in a functional way; therefore, learning of the state flow would become an easier task. Note that to ensure $F_{\boldsymbol{\phi}}(\mathbf{x}_{0},0)=R(\mathbf{x}_{0})$, we need to force $\tilde{F}_{\boldsymbol{\phi}}(\mathbf{x}_{0},0)=1$ for all $\mathbf{x}_{0}$ at the terminal step.
Incorporating denoising diffusion-specific structure
However, the intermediate state $\mathbf{x}_{t}$ is noisy under our context, and thus not appropriate for being evaluated by the given reward function, which would give noisy result. What’s more, what we are interested here is to “foresee” the reward of the terminal state $\mathbf{x}_{0}$ taken from the (partial) trajectory $\mathbf{x}_{t:0}$ starting from given $\mathbf{x}_{t}$. As a result, we can do the FL technique utilizing the particular structure of diffusion model as in $F_{\boldsymbol{\phi}}(\mathbf{x}_{t},t)=\tilde{F}_{\boldsymbol{\phi}}(\mathbf{x}_{t},t)R(\hat{\mathbf{x}}_{\boldsymbol{\theta}}(\mathbf{x}_{t},t))$, where $\hat{\mathbf{x}}_{\boldsymbol{\theta}}$ is the data prediction network. We notice that a similar technique has been used to improve classifier guidance (Bansal2023UniversalGF). In short, our innovation in FL technique is
$F_{\boldsymbol{\phi}}(\mathbf{x}_{t},t)=\tilde{F}_{\boldsymbol{\phi}}(\mathbf{x}_{t},t)R(\mathbf{x}_{t})\ \Longrightarrow$ (3)
$F_{\boldsymbol{\phi}}(\mathbf{x}_{t},t)=\tilde{F}_{\boldsymbol{\phi}}(\mathbf{x}_{t},t)R(\hat{\mathbf{x}}_{\boldsymbol{\theta}}(\mathbf{x}_{t},t)).$ (4)
Then the FL-DB training objective $\ell_{\text{FL}}(\mathbf{x}_{t},\mathbf{x}_{t-1})$ becomes
$\left(\log\frac{\tilde{F}_{\boldsymbol{\phi}}(\mathbf{x}_{t},t)R(\hat{\mathbf{x}}_{\boldsymbol{\theta}}(\mathbf{x}_{t},t))p_{\boldsymbol{\theta}}(\mathbf{x}_{t-1}|\mathbf{x}_{t})}{\tilde{F}_{\boldsymbol{\phi}}(\mathbf{x}_{t-1},{t-1})R(\hat{\mathbf{x}}_{\boldsymbol{\theta}}(\mathbf{x}_{t-1},{t-1}))q(\mathbf{x}_{t}|\mathbf{x}_{t-1})}\right)^{2}.$ (5)
Remark 1 (GPU memory and the choice of GFlowNet objectives).
Similar to the temporal difference-$\lambda$ in RL (sutton1988learning), it is possible to use multiple connected transition steps rather than a single transition step to construct the learning objective. Other GFlowNet objectives such as malkin2022trajectory; madan2022learning use partial trajectories with a series of transition steps to construct the training loss and provide a different trade-off between variance and bias in credit assignment. However, for large-scale setups, this is not easy to implement, as computing policy probabilities for multiple transitions would correspondingly increase the GPU memory and computation multiple times. For example, in the Stable Diffusion setting, we could only use a batch size of $8$ on each GPU for single transition computation. If we want to use a two transition based training loss, we would need to decrease the batch size by half to $4$. Similarly, we will have to shorten the trajectory length by a large margin if we want to use trajectory balance. This may influence the image generation quality and also make it tricky to compare with the RL baseline, which can be implemented with single transitions and does not need to decrease batch size or increase gradient accumulation. In practice, we find that single transition algorithms (such as our RL baseline) perform reasonably well.
2.2 A KL-based GFlowNet algorithm with REINFORCE gradient
$\vspace{-0.2cm}\min_{\boldsymbol{\theta}}\mathcal{D}_{\mathrm{KL}}\left(p_{\boldsymbol{\theta}}(\mathbf{x}_{t-1}|\mathbf{x}_{t})\|\frac{F_{\boldsymbol{\phi}}(\mathbf{x}_{t-1},{t-1})q(\mathbf{x}_{t}|\mathbf{x}_{t-1})}{F_{\boldsymbol{\phi}}(\mathbf{x}_{t},t)}\right).\vspace{-0.2cm}$ (6)
In theory, when DB is perfectly satisfied, the right term $F_{\boldsymbol{\phi}}(\mathbf{x}_{t-1},{t-1})q(\mathbf{x}_{t}|\mathbf{x}_{t-1})/{F_{\boldsymbol{\phi}}(\mathbf{x}_{t},t)}$ is a normalized density; in practice, it could be an unnormalized one but does not affect the optimization. Next, define $b(\mathbf{x}_{t},\mathbf{x}_{t-1})$ to be
[…]
Proposition 2.
$\nabla_{\boldsymbol{\theta}}\mathcal{D}_{\mathrm{KL}}\left(p_{\boldsymbol{\theta}}(\mathbf{x}_{t-1}|\mathbf{x}_{t})\|\frac{F_{\boldsymbol{\phi}}(\mathbf{x}_{t-1},{t-1})q(\mathbf{x}_{t}|\mathbf{x}_{t-1})}{F_{\boldsymbol{\phi}}(\mathbf{x}_{t},t)}\right)$ (8)
$=\mathbb{E}_{\mathbf{x}_{t-1}\sim p_{\boldsymbol{\theta}}(\cdot|\mathbf{x}_{t})}\left[b(\mathbf{x}_{t},\mathbf{x}_{t-1})\nabla_{\boldsymbol{\theta}}\log p_{\boldsymbol{\theta}}(\mathbf{x}_{t-1}|\mathbf{x}_{t})\right].$ (9)
We defer its proof to Section C.1 and make the following remarks:.
Remark 3 (gradient equivalence to detailed balance).
Therefore, we can define a new objective $\ell_{\text{KL}}(\mathbf{x}_{t},\mathbf{x}_{t-1})$
$b(\mathbf{x}_{t},\mathbf{x}_{t-1})\ \text{clip}\left(\frac{p_{\boldsymbol{\theta}}(\mathbf{x}_{t-1}|\mathbf{x}_{t})}{p_{{\boldsymbol{\theta}}_{\text{old}}}(\mathbf{x}_{t-1}|\mathbf{x}_{t})},1-\epsilon,1+\epsilon\right),$ (11)
where $\mathbf{x}_{t-1}\sim p_{{\boldsymbol{\theta}}_{\text{old}}}(\cdot|\mathbf{x}_{t})$. Here we also introduce a clip operation to remove too drastic update, following PPO (schulman2017proximal). We use this to update the policy parameter ${\boldsymbol{\theta}}$ and use FL-DB to only update ${\boldsymbol{\phi}}$. We call this “diffusion alignment with GFlowNet and REINFORCE gradient” method to be DAG-KL. Note that when calculating $b(\mathbf{x}_{t},\mathbf{x}_{t-1})$, we also adopt the diffusion-specific FL technique developed in Section 2.1. We also put the algorithmic pipeline of DAG-KL in Algorithm 1.
3 Experiments
Experimental setups
We choose Stable Diffusion v1.5 (Rombach2021HighResolutionIS) as our base generative model. For training, we use low-rank adaptation (hu2021lora, LoRA) for parameter efficient computation. As for the reward functions, we do experiments with the LAION Aesthetics predictor, a neural aesthetic score trained from human feedback to give an input image an aesthetic rating. For text-image alignment rewards, we choose ImageReward (Xu2023ImageRewardLA) and human preference score (HPSv2) (Wu2023HumanPS). They are both CLIP (Radford2021LearningTV)-type models, taking a text-image pair as input and output a scalar score about to what extent the image follows the text description. We also test with the (in)compressibility reward, which computes the file size if the input image is stored in hardware storage. As for the prompt distribution, we use a set of $45$ simple animal prompts from Black2023TrainingDM for the Aesthetics task; we use the whole imagenet classes for the (in)compressibility task; we use the DrawBench (saharia2022photorealistic) prompt set for the ImageReward task; we use the photo and painting prompts from the human preference dataset (HPDv2) (Wu2023HumanPS) for the HPSv2 task. We notice that in our experiments, we use prompt set containing hundreds of prompts which is more than some previous work such as Black2023TrainingDM.
Effectiveness of the proposed methods
We first demonstrate that our proposed methods could generated images that have meaningful improvements corresponding to the rewards being used. In Figure 1, we compare the images from the original Stable Diffusion pretrained model and our proposed method. After our post-training, the generated images become more vibrant and vivid; we also notice that these images have slightly higher saturation, which we believe is aligned with the human preference on good-looking pictures. We also visualize the experiment results on compressibility and incompressibility tasks in Figure 3. The second row shows the generated images from the model trained with the compressibility reward, which have low details and smooth textures, and also have very limited colors. On the other hand, the model trained with incompressibility reward would generate images with high frequency texture, as shown in the third row. These results indicate that our method could effectively incorporate the reward characteristics into the generative models. We defer more experimental details to Section D.2.
Algorithmic comparisons
The main baseline we compare with is denoising diffusion policy optimization (Black2023TrainingDM, DDPO), an RL algorithm that is specifically designed for denoising diffusion alignment and has been shown to outperform other align-from-black-box-reward methods including (Lee2023AligningTM; Fan2023DPOKRL). We show the reward curves w.r.t. the training steps of the aesthetic, ImageReward, and HPSv2 rewards in Figure 2. Here, the number of training steps corresponds proportionally to the number of trajectories collected. Both our proposed methods, DAG-DB and DAG-KL, achieve faster credit assignment than the DDPO baseline by a large margin. We also put corresponding curve plots for compressibility and incompressibility rewards in Figure 6, which also demonstrates the advantage of our methods. We defer related training details to Section D.2.
Appendix A Preliminaries
A.1 Diffusion models
@@
Denoising diffusion model (Vincent2011ACB; SohlDickstein2015DeepUL; Ho2020DenoisingDP; Song2020ScoreBasedGM) is a class of hierarchical latent variable models. The latent variables are initialized from a white noise $\mathbf{x}_{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ and then go through a sequential denoising (reverse) process $p_{\boldsymbol{\theta}}(\mathbf{x}_{t-1}|\mathbf{x}_{t})$. Therefore, the resulting generated distribution takes the form of
@@
On the other hand, the variational posterior $q(\mathbf{x}_{1:T}|\mathbf{x}_{0})$, also called a diffusion or forward process, can be factorized as a Markov chain $\prod_{t=1}^{T}q(\mathbf{x}_{t}|\mathbf{x}_{t-1})$ composed by a series of conditional Gaussian distributions $q(\mathbf{x}_{t}|\mathbf{x}_{t-1})=\mathcal{N}(\mathbf{x}_{t};{\alpha_{t}/\alpha_{t-1}}\mathbf{x}_{t-1},(1-\alpha_{t}^{2}/\alpha_{t-1}^{2})\mathbf{I})$, where $\{\alpha_{t},\sigma_{t}\}_{t}$ is a set of pre-defined signal-noise schedule. Specifically, in Ho2020DenoisingDP we have $\alpha_{t}^{2}+\sigma_{t}^{2}=1$. The benefit of such a noising process is that its marginal has a simple close form: $q(\mathbf{x}_{t}|\mathbf{x}_{0})=\int q(\mathbf{x}_{1:t}|\mathbf{x}_{0})\mathop{}\!\mathrm{d}\mathbf{x}_{1:t-1}=\mathcal{N}(\mathbf{x}_{t};\alpha_{t}\mathbf{x}_{0},\sigma_{t}^{2}\mathbf{I})$.
@@
where $\hat{\mathbf{x}}_{\boldsymbol{\theta}}(\mathbf{x}_{t},t)$ is a deep neural network to predict the original clean data $\mathbf{x}_{0}$ given the noisy input $\mathbf{x}_{t}=\alpha_{t}\mathbf{x}_{0}+\sigma_{t}{\boldsymbol{\epsilon}}$, which can be used to parameterize the denoising process $p_{\boldsymbol{\theta}}(\mathbf{x}_{t-1}|\mathbf{x}_{t})=\mathcal{N}(\mathbf{x}_{t-1};\left(\sigma_{t-1}^{2}\alpha_{t}\mathbf{x}_{t}+(\alpha_{t-1}^{2}-\alpha_{t}^{2})\hat{\mathbf{x}}_{\boldsymbol{\theta}}(\mathbf{x}_{t},t)\right)/\sigma_{t}^{2}\alpha_{t-1},(1-\alpha_{t}^{2}/\alpha_{t-1}^{2})\mathbf{I})$. In practice, the network can also be parameterized with noise prediction or v-prediction (salimans2022progressive). The network architecture usually has a U-Net (Ronneberger2015UNetCN) structure.
@@
A.2 GFlowNets
Generative flow network (bengio2021flow, GFlowNet) is a high-level algorithmic framework of amortized inference, also known as training generative models with a given unnormalized target density function. Let $\mathcal{G}=(\mathcal{S},\mathcal{A})$ be a directed acyclic graph, where $\mathcal{S}$ is the set of states and $\mathcal{A}\subseteq\mathcal{S}\times\mathcal{S}$ are the set of actions. We assume the environmental transition is deterministic, i.e., one action would only lead to one next state. There is a unique initial state $\mathbf{s}_{0}\in\mathcal{S}$ which has no incoming edges and a set of terminal states $\mathbf{s}_{N}$ without outgoing edges. A GFlowNet has a stochastic forward policy $P_{F}(\mathbf{s}^{\prime}|\mathbf{s})$ for transition $(\mathbf{s}\to\mathbf{s}^{\prime})$ as a conditional distribution over the children of a given state $\mathbf{s}$, which can be used to induce a distribution over trajectories via $P(\tau)=\prod_{n=0}^{N-1}P_{F}(\mathbf{s}_{n+1}|\mathbf{s}_{n})$, where $\tau=(\mathbf{s}_{0},\mathbf{s}_{1},\ldots,\mathbf{s}_{N})$. On the other hand, the backward policy $P_{B}(\mathbf{s}|\mathbf{s}^{\prime})$ is a distribution over the parents of a given state $\mathbf{s}^{\prime}$. The terminating distribution defined by $P_{T}(\mathbf{x})=\sum_{\tau\to\mathbf{x}}P_{F}(\tau)$ is the ultimate terminal state distribution generated by the GFlowNet. The goal of training GFlowNet is to obtain a forward policy such that $P_{T}(\cdot)\propto R(\cdot)$, where $R(\cdot)$ is a black-box reward function or unnormalized density that takes only non-negative values. Notice that we do not know the normalizing factor $Z=\sum_{\mathbf{x}}R(\mathbf{x})$. We can use the trajectory flow function $F(\tau)=ZP_{F}(\tau)$ to take in the effect of the normalizing factor, and the corresponding state flow function $F(\mathbf{s})=\sum_{\tau\ni\mathbf{s}}F(\tau)$ to model the unnormalized probability flow of intermediate state $\mathbf{s}$.